A ball of mass 100g is thrown vertically upwards from a point 2m above ground level with a speed of 14m/s.
a. With an origin at ground level, find the total mechanical energy of the ball when it is travelling at speed v m/s at a heigh h m.
b. Assuming that mechanical energy is conserved show that v^2+20h=236
c. Calculate the greatest height reached by the ball
d. Calculate the speed with which the ball hits the ground
1
Expert's answer
2020-03-13T10:53:39-0400

a)

"E=0.5mv^2+mgh"

b)


"0.5mv^2+m(10)h=0.5m(14)^2+m(10)(2)"

"v^2+20h=(14)^2+2(10)(2)=236"

c) It will be when speed is zero:


"20H=236"

"H=11.8\ m"

d)It will be when height is zero:


"v^2=236"

"v=15.4\frac{m}{s}"